Calculate the value for the mass defect of a mole of the following:



a. Boron – 15 (isotopic mass = 15.031088 u)



b. Carbon – 14 (isotopic mass = 14.003241 u)



c. Nitrogen – 17 (isotopic mass = 17.008449 u)

1
Expert's answer
2022-06-15T17:18:03-0400

1 u = 1.661*10-24 g

According to Avogadro's law, 1 mole of a substance contains 6.02 * 1023 particles (atoms, molecules).

To find the mass of 1 mole of isotope, multiply the given mass by 1.661 * 10-24 and then multiply by the avogadro's number (6.02 * 1023).


a.)

M = 15.031088 * 1.661*10-24 * 6.02*1023 = = 15 g



b.)

M = 14.003241 * 1.661*10-24 * 6.02*1023 = = 14 g



c.)

M = 17.008449 * 1.661*10-24 * 6.02*1023 =

= 17 g